OpenStudy (anonymous):

How do you integrate 2sin5xsinx?

OpenStudy (mathmale):

I did a search for "trig sum and difference identities" and came up with the following URL: http://www.purplemath.com/modules/idents.htm Within that page, look for the heading "Sum Identities." Hope this gets you onto the path towards solving this problem. Hint: Suppose a+b=5 and a-b =1 Solve for a and b.

OpenStudy (anonymous):

So 2a=6 a=3 3+b=5 b=2 Then what do I do next?

OpenStudy (anonymous):

Also, does the formula under "Product Identities" work as well?

OpenStudy (mathmale):

Actually, those formulas under Product identities are faster to use. Glad that you've noticed that. Starting with 2sin5xsinx, I'd hold the factor, 2, and concentrate on the sin5x*sinx part. Looks like the very last one, for sin x*sin y, is the one we want.

OpenStudy (mathmale):

Care to try evaluating sin 5x * sin x using this formula? When done, be certain to multiply the entire result by 2. (Why?)

OpenStudy (anonymous):

So using sin(P)sin(Q)=1/2[ 2cos(x-y) - 2cos(x+y) ] = 1/2 [ 2cos (5x-x) - 2cos (5x+x) ] = 1/2 [ 2cos4x - 2cos6x] Would the final answer be this - = (1/2) (1/4) 2sin4x - (1/2) (1/6) 2sin6x = 1/4 sin4x - 1/6 sin6x ???